OpenStudy (anonymous):

Could somebody show me the steps to simplifying this equation? (n+5)(n+4)!

OpenStudy (anonymous):

it will be (n+5)!

OpenStudy (anonymous):

how do you know that and how did you figure it out? @madrockz

OpenStudy (anonymous):

(n+5)! will be (n+5)(n+4)(n+3).... or equal to (n+5)(n+4)!. {(n+4)!=(n+4)(n+3)..} The factorial of a number can always be expressed as the product of an integer and a factorial of a number smaller than it

OpenStudy (akashdeepdeb):

@BaileyC Di you get it? Or do you want me to explain?

OpenStudy (anonymous):

okay, so it is (n+5)! because it is equal to (n+4)! because its a factorial?... @madrockz @AkashdeepDeb

OpenStudy (anonymous):

See, if someone could show me the solving of this problem step by step, that would be suuuper helpful

OpenStudy (akashdeepdeb):

Not quite. Let us say we had 3 times 2! Now that would be 3! [Because 3 times 2! = 3*2*1 = 3! ] Similarly, (n+5)[ (n+4)! ] = (n+5)[(n+4)(n+3)(n+2).........3.2.1 ] Which actually is = (n+5)! Understood? :)

OpenStudy (anonymous):

okay. thank you.

OpenStudy (akashdeepdeb):

:)